Tom is three times as old as Mark. Twice Mark's age plus Tom's age is 90. How old is each person
Found 2 solutions by stanbon, rwm:
Answer by stanbon(75887)

Tom is three times as old as Mark. Twice Mark's age plus Tom's age is 90. How old is each person
---
Equations:
t = 3m
2m + t = 90
----
Substitute for "t" and solve for "m":
2m + 3m = 90
5m = 90
m 18 (Mark's age now)
t = 3m = 54 (Tom's age now)
===============================
Cheers,
Stan H.
=============

Answer by rwm(914)

we have two unknowns so we need two equations.
t=3m
2m+t=90
2m+3m=90
5m=90
m=18
t=54
check
36+54=90
ok